# Conditional Density Estimations from Privacy-Protected Data

### v3

@@ 3.1 Private data likelihood estimation
Our first strategy is to approximate the private data marginal likelihood $f(s_{\mathrm{dp}}\mid\theta)$ with a neural likelihood estimator $q_{\phi}(s_{\mathrm{dp}}\mid\theta)$. When training $q_{\phi}(s_{\mathrm{dp}}\mid\theta)\approx f(s_{\mathrm{dp}}\mid\theta)$, we aim to minimize their average KL divergence under the prior $\pi(\theta)$, corresponding to minimizing
@@
After some derivations (in Section S1), Eq.(4) is equivalent to $\mathbb{E}_{\theta,s_{\mathrm{dp}}}\left[-\log q_{\phi}(s_{\mathrm{dp}}\mid\theta)\right]$ up to a constant independent of $\phi$, where the expectation is taken with respect to $(\theta,s_{\mathrm{dp}})\sim\pi(\theta)\cdot f(s_{\mathrm{dp}}\mid\theta)$. Alternatively, we can integrate with respect to the confidential data generating process, with the loss function
@@
The resulting private data likelihood estimation is $\widehat{\phi}=\arg\min\ell_{\mathrm{PLE}}(\phi)$ and $q_{\widehat{\phi}}(s_{\mathrm{dp}}\mid\theta)$.
When the primary inference goal is the maximum likelihood estimator, one can approximate it with $\hat{\theta}_{\textnormal{MLE}}=\arg\max_{\theta}q_{\phi}(s_{\mathrm{dp}}\mid\theta).$ Under the Bayesian paradigm, the posterior approximation of Eq.(3) can be $\widehat{\pi}_{\textnormal{PLE}}(\theta)\propto\pi(\theta)q_{\widehat{\phi}}(s_{\mathrm{dp}}\mid\theta)$. Quantities such as posterior median, mean, and credible regions can be estimated accordingly.
@@ 3.2 Private data posterior estimation
We can also approximate the private data posterior $\pi(\theta\mid s_{\mathrm{dp}})$ in Eq.(3) directly, bypassing the synthetic likelihood step, with a neural posterior estimator $q_{\phi}(\theta\mid s_{\mathrm{dp}})$. Learning $q_{\phi}(\theta\mid s_{\mathrm{dp}})\approx\pi(\theta\mid s_{\mathrm{dp}})$, we can minimize their KL divergence.
$\mathcal{D}_{\mathrm{KL}}\left(\pi(\theta\mid s_{\mathrm{dp}})\|q_{\phi}(\theta\mid s_{\mathrm{dp}})\right)=\mathbb{E}_{\pi(\theta\mid s_{\mathrm{dp}})}\left[\log\frac{\pi(\theta\mid s_{\mathrm{dp}})}{q_{\phi}(\theta\mid s_{\mathrm{dp}})}\right].$ (6)
With derivations in Section S1, we can write the loss function as
@@
Many Bayesian problems assume an uninformative prior $\pi(\theta)$, which are dispersed in the parameter space instead of concentrating on high density regions of the posterior. As a result, a naive Monte Carlo strategy to approximate the expectation in Eq.(7) will be inefficient. Similar to importance sampling, where one utilize a proposal distribution $\tilde{p}(\theta)$ to reduce variance during numerical integration, we can rewrite Eq.(7) as an expectation with respect to some proposal distribution $\tilde{p}$. Greenberg et al., (2019) proposed the Automatic Posterior Transformation (APT) method, with
$\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}}):=q_{\phi}(\theta\mid s_{\mathrm{dp}})\frac{\tilde{p}(\theta)}{\pi(\theta)}\frac{1}{Z(s_{\mathrm{dp}},\phi)},$
where $Z(s_{\mathrm{dp}},\phi)$ is a normalization constant and can be estimated as $\sum_{\theta^{\prime}\in\Theta}q_{\phi}(\theta^{\prime}\mid s_{\mathrm{dp}})/\pi(\theta^{\prime})$, $\theta^{\prime}\sim\tilde{p}(\theta)$ using importance sampling. In this case, the loss function based on APT is
$\ell_{\mathrm{PPE-A}}(\phi)=\mathbb{E}_{\tilde{p}(\theta,x)}\left[-\int_{\mathbb{S}}\eta(s_{\mathrm{dp}}\mid x)\log\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}})\mathrm{d}s_{\mathrm{dp}}\right].$ (8)
This APT-based objective function will be useful for the sequential training strategy we introduce in Section 3.4. The derivation of Eq.(8) is in Section S1.
@@ 3.3 Nested RQMC estimators
Both loss functions in Eq.(5) and (8) are double integrals, where the outer integral is an expectation with respect to some proposal distribution $\tilde{p}$ and the inner integral involves the privacy mechanism $\eta$ and the neural estimator $q_{\phi}$. We can inspect their general forms
@@
Here $g(s_{\mathrm{dp}},\theta)=\log\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}})$ for PPE-A and $g(s_{\mathrm{dp}},\theta)=\log q_{\phi}(s_{\mathrm{dp}}\mid\theta)$ for PLE.
With iid samples from the joint density $\{(\theta^{(i)},x^{(i)})\}_{i=1}^{N}\sim\tilde{p}(\theta,x)=\tilde{p}(\theta)f(x\mid\theta)$, we can unbiasedly approximate Eq.(9) with
@@
The inner integrals $I(\theta,x)=\int\eta(s_{\mathrm{dp}}\mid x)g(s_{\mathrm{dp}},\theta)\mathrm{d}s_{\mathrm{dp}}$ can be approximated with standard Monte Carlo integration techniques, as popular DP mechanisms such as Laplace, Gaussian, and Exponential can be easily simulated. For MC strategies with $M$ samples, the root-mean-squared-error (RMSE) of $I(\theta,x)$ approximations are typically on the order of $\mathcal{O}(1/\sqrt{M})$.
In many applications, the DP query result $s_{\mathrm{dp}}$ is a private descriptive statistic of $x$, which means $s_{\mathrm{dp}}$ is embedded in a low-dimensional space $\mathbb{S}$ with $r=\mathrm{dim}(\mathbb{S})\leq\mathrm{dim}(\mathbb{X}^{n}).$ Let $u\sim\mathcal{U}[0,1]^{r}$ be a random number from the $r$-dimensional hypercube, fixing $x$, we can write the mechanism by $\tau:[0,1]^{r}\to\mathbb{S}$ with $\tau(u;s(x))$.
We can leverage randomized quasi-Monte Carlo (RQMC) (Owen,, 1997; L’Ecuyer,, 2018) to efficiently integrate over low-dimensional spaces and achieve favorable convergence rates for numerical integration. RQMC generates a low-discrepancy sequence $\{v^{(1)},\cdots,v^{(M)}\}\subset[0,1]^{r}$ to achieve the estimator
@@
In addition to being unbiased, the RQMC estimator also has RMSE compared to MC estimators. The $\star$-discrepancy of the point set $\{v^{(1:M)}\}$, denoted by $D^{\star}(v^{(1:M)})$ is of the order $\mathcal{O}(M^{-1}(\log M)^{r})=\mathcal{O}(M^{-1+\delta})$ for a positive constant $\delta$. If our neural approximation family $\tilde{g}_{\theta,x}(\cdot)$ has bounded Hardy-Krause variation $V_{\mathrm{HK}}[\tilde{g}_{\theta,x}(\cdot)]$ (Aistleitner et al.,, 2017), then, according to Basu and Owen, (2016), the variance of the RQMC estimator in Eq.(11) satisfies
$\mathrm{Var}(\hat{I}_{\theta,x}^{\mathrm{RQMC}})\leq V^{2}_{\mathrm{HK}}(\tilde{g}_{\theta,x})(D^{\star})^{2}=\mathcal{O}(M^{-2+\delta}).$ (12)
Thus, the RMSE of RQMC is of the order $\mathcal{O}(M^{-1+\delta/2)}$, achieving faster convergence than an MC estimator. We verify this improvement in convergence from the RQMC estimator on the SIR model and linear regression example with the neural spline flow approximation family (Durkan et al.,, 2019), in Fig. 1.
Figure 1: Rate of convergence of MC and RQMC.
@@ 3.4 Sequential neural estimations
This section discusses our sequential neural approximation methods on privacy-protected data. We summarize the Sequential Private Posterior Estimation (SPPE) algorithm in Algo.1 and the Sequential Private Likelihood Estimation (SPLE) in Algo.2.
@@
     Update $\phi\leftarrow\arg\min_{\phi}\hat{\ell}_{\mathrm{PPE-A}}(\phi)$ using $\hat{I}^{\mathrm{RQMC}}$ (11) on $\mathcal{D}_{r}$
@@
return $\hat{\pi}(\theta\mid s_{\mathrm{dp}}^{o})=\tilde{p}_{R}(\theta)$
@@
     Update $\phi\leftarrow\mathop{\arg\min}_{\phi}\hat{\ell}_{\mathrm{PLE}}(\phi)$ using $\hat{I}^{\mathrm{RQMC}}$ (11) on $\mathcal{D}_{r}$
@@
Our methods refine the normalizing flows $q_{\phi}$ to approximate the target distributions by iteratively minimizing the loss functions taking the general form in Eq.(9). In Eq.(9), using a proposal distribution $\tilde{p}(\theta,x)$ that overlaps more with the joint posterior distribution $\pi(\theta,x\mid s_{\mathrm{dp}})$ is more efficient than using simulations from the joint distribution $\pi(\theta)f(x\mid\theta)$, in terms of variance reduction for Monte Carlo integration. To this end, we sequentially train the neural density estimators: after round $r$, the posterior approximation $\widehat{q}^{(r)}(\theta)\approx\pi(\theta\mid s_{\mathrm{dp}})$ becomes a proposal distribution in the loss functions $\ell(\phi)$ for round $(r+1).$ This approach can gradually move $q_{\phi}$ towards high density regions of the private data posterior, and thus achieve good accuracy with fewer samples from the confidential data simulator.
@@ 4.1 SIR model for disease spread
The SIR model is a time-series model that describes how an infectious disease spread in a closed population. It is is most often used as a deterministic ordinary differential equation (ODE), but can also be represented by a Markov jump process.
@@
Privatizing the infection curve. We propose a mechanism to publish privatized data about the infection process. This algorithm adds calibrated noise to the SIR process, protecting the individual’s infection status while allowing analysts to make inferences about population parameters $(\beta,\gamma)$. At each time, the proportion of infected individuals is $I(t)/K$. We shall privatize this infection trajectory.
Proposition 4.
@@
Consider a sequence of $L$ points in the time interval $[0,T]$, our privatized query can be $s_{\mathrm{dp}}=(s_{1},\ldots,s_{L})$ where each $s_{i}\sim\mathrm{Binomial}\left(n,\frac{I(t_{i})+m}{K+2m}\right)$ independently. The mechanism generating $s_{\mathrm{dp}}=(s_{1},\ldots,s_{L})$ satisfies $\epsilon$-DP, with $\epsilon=\frac{n}{m}L.$
Here $s_{\mathrm{dp}}$ is a differential private time series, reflecting the speed of disease spread in a population.
@@
Experiments on synthetic data. We illustrate the performance of SPPE and SPLE on synthetic privatized SIR model data. The data generating parameters are set to emulate a measles outbreak. We describe prior specifications and implementation details in the Appendix.
@@
Fig. 2 describes the convergence of the posterior approximations $\hat{\pi}(\theta\mid s_{\mathrm{dp}}^{o})$ and compares our results with SMC-ABC (Beaumont et al.,, 2009). SMC-ABC is chosen as the baseline comparison as it aims to recover the posterior exactly without resorting to variational approximations. We run the SMC-ABC method with up to $5\times 10^{5}$ simulations. Both SPPE and SPLE quickly adapt to meet the SMC-ABC results, with order of magnitudes fewer simulations needed. After the first round of simulations, SPPE can identify the high probability region of $\pi(\theta\mid s_{\mathrm{dp}})$, while the SPLE posterior is still exploring the parameter space. See SI for the SPLE approximation after $r=1.$ After $r=5$ rounds, both methods can concentrate around the posterior mean and have captured the posterior correlation $\mathrm{cov}(\beta,\gamma\mid s_{\mathrm{dp}}).$
Figure 2: Convergence of sequential posterior estimations given DP-protected infection trajectory under the SIR model. Each round entails $N=1000$ simulations. Orange: SPPE; red: SPLE; gray: SMC-ABC.
By inspecting marginal posterior histograms (See Appendix), we find that, in this example, the posterior approximated by SPPE is slightly more concentrated than those from SMC-ABC and SPLE.
@@
To evaluate the performance of our methods, we use the following metrics: (1) MMD (Gretton et al.,, 2012): maximum mean discrepancy between the neural estimated posteriors and SMC-ABC posterior; (2) C2ST (Lopez-Paz and Oquab,, 2016): classifier two sample tests; (3) NLOG: negative log density at true parameters; and (4) LMD: log median distance from simulated to observed $s_{\mathrm{dp}}.$ Smaller values indicate better performance for all four metrics. In Fig. 3, we compare the performance of these methods at various numbers of simulation samples/rounds. After Round 5, SPLE and SPPE have similar accuracy. See Table S1 in the Appendix for run-time comparisons of SPPE and SPLE against SMC-ABC.
@@
Figure 3: Approximation accuracy by SPPE (orange) and SPLE (red) on the SIR model against number of rounds, the error bars represent the mean with the upper and lower quartiles.
@@
Figure 4: Inference on real infectious disease outbreaks. A. Visualization of the posterior distribution given private infection curve on applied to flu, Ebola [in a) Guinea, b) Liberia, and c) Sierra Leone], and Covid-19 in Clark County, Nevada. datasets. B. Mean and 95% credible intervals for $R_{0}=\beta/\gamma$ with different methods in each dataset. Grey: SMC-ABC; orange: SPPE; red: SPLE.
We apply our privacy mechanism and inference methods (SPPE and SPLE) on several real infectious disease outbreaks: influenza, Ebola, and Covid-19. In Fig. 4-A, we compare the posterior distributions $\pi(\beta,\gamma\mid s_{\mathrm{dp}})$ obtained from SMC-ABC, SPPE, and SPLE. The results are similar to synthetic data experiments: when SPPE and SPLE use the same computational resources, the SPPE posterior converges faster than SPLE. In Fig. 4-B, we inspect the 95%-credible interval for the ratio $R_{0}=\beta/\gamma$, which is known as the basic reproduction number. Our $R_{0}$ estimates using privatized data are consistent with common estimates of $R_{0}$ for these diseases (Eisenberg,, 2020), with the exception of the flu outbreak (which should be modeled by the SI model instead of SIR).

### v4

@@ 3.2 Private data likelihood estimation
Our first neural density approximation strategy is to approximate the private data marginal likelihood $f(s_{\mathrm{dp}}\mid\theta)$ with a neural network denoted by $q_{\phi}(s_{\mathrm{dp}}\mid\theta)$. When training $q_{\phi}(s_{\mathrm{dp}}\mid\theta)\approx f(s_{\mathrm{dp}}\mid\theta)$, we aim to minimize their average KL divergence under the prior $\pi(\theta)$, corresponding to minimizing
@@
After some derivations (in Appendix A), Equation (4) is equivalent to $\mathbb{E}_{p(\theta,s_{\mathrm{dp}})}\left[-\log q_{\phi}(s_{\mathrm{dp}}\mid\theta)\right]$ up to a constant independent of $\phi$. The expectation is taken with respect to the prior and the private data generating process $p(\theta,s_{\mathrm{dp}})=\pi(\theta)\cdot f(s_{\mathrm{dp}}\mid\theta)$, which is intractble. To facilitate computations, let’s write Equation (4) with respect to the confidential data generating process, resulting in
@@
With $\widehat{\phi}=\arg\min\ell_{\mathrm{PLE}}(\phi)$, the resulting private data likelihood estimation is $q_{\widehat{\phi}}(s_{\mathrm{dp}}\mid\theta)$. Since Equation (5) is an expectation with respect to a tractable distribution, we can approximate the integral with Monte Carlo methods, discussed in Section 3.4.
When the primary inference goal is the maximum likelihood estimator, one can approximate it with $\widehat{\theta}_{\textnormal{MLE}}=\arg\max_{\theta}q_{\widehat{\phi}}(s_{\mathrm{dp}}\mid\theta).$ Under the Bayesian paradigm, the posterior approximation of Equation (3) can be $\widehat{\pi}_{\textnormal{PLE}}(\theta)\propto\pi(\theta)q_{\widehat{\phi}}(s_{\mathrm{dp}}\mid\theta)$. Quantities such as posterior median, mean, and credible regions can be estimated accordingly.
@@ 3.3 Private data posterior estimation
Now we approximate the private data posterior $\pi(\theta\mid s_{\mathrm{dp}})$ in Equation (3) directly with a neural posterior estimator $q_{\phi}(\theta\mid s_{\mathrm{dp}})$, bypassing the synthetic likelihood step in Section 3.2. To find $q_{\phi}(\theta\mid s_{\mathrm{dp}})\approx\pi(\theta\mid s_{\mathrm{dp}})$, let us minimize their KL divergence
$\mathcal{D}_{\mathrm{KL}}\left(\pi(\theta\mid s_{\mathrm{dp}})\|q_{\phi}(\theta\mid s_{\mathrm{dp}})\right)=\mathbb{E}_{\pi(\theta\mid s_{\mathrm{dp}})}\left[\log\frac{\pi(\theta\mid s_{\mathrm{dp}})}{q_{\phi}(\theta\mid s_{\mathrm{dp}})}\right].$
As shown in Section A, this is equivalent to
@@
Many Bayesian problems use uninformative priors $\pi(\theta)$, which are dispersed in the parameter space. As a result, when employing a naive Monte Carlo strategy that uses the prior as the proposal to approximate expectations in Equation (5) or Equation (6) has most of its samples falling in low-density regions, leading to inefficient estimation due to high variance. To address this issue, importance sampling utilizes a proposal distribution $\tilde{p}$ to change the bases of integration and thus reduce the variance of numerical integration. The automatic posterior transformation (APT) method (Greenberg et al.,, 2019) uses $\tilde{p}(\theta,x)=\tilde{p}(\theta)f(x\mid\theta)$ as proposal and has (unnormalized) importance weights
$\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}})\propto q_{\phi}(\theta\mid s_{\mathrm{dp}})\frac{\tilde{p}(\theta)}{\pi(\theta)}.$ (7)
Based on this idea, we design a modified loss of Equation (6) that leverages the APT framework to improve efficiency in posterior estimation. Specifically, we propose the following loss function
$\ell_{\mathrm{PPE-A}}(\phi)=\mathbb{E}_{\tilde{p}(\theta,x)}\left[-\int_{\mathbb{S}}\eta(s_{\mathrm{dp}}\mid x)\log\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}})\mathrm{d}s_{\mathrm{dp}}\right],$ (8)
and it is useful for the sequential training strategy we introduce in Section 3.5. The derivation of Equation (8) is in Appendix A.
@@ 3.4 Nested RQMC estimators
We can inspect the general form of loss functions in Equations (5) and (8) with
@@
Here $g(s_{\mathrm{dp}},\theta)=\log\tilde{q}_{\phi}(\theta\mid s_{\mathrm{dp}})$ for PPE-A and $g(s_{\mathrm{dp}},\theta)=\log q_{\phi}(s_{\mathrm{dp}}\mid\theta)$ for PLE. Equation 9 is a double integral, where the outer expectation is with respect to some proposal distribution $\tilde{p}$ and the inner integral involves the privacy mechanism $\eta$ and the neural estimator $q_{\phi}$.
With independent and identically distributed (i.i.d.) samples from the joint density $\{(\theta^{(i)},x^{(i)})\}_{i=1}^{N}\sim\tilde{p}(\theta,x)=\tilde{p}(\theta)f(x\mid\theta)$, we can unbiasedly approximate Equation (9) with
@@
The inner integrals $I(\theta,x)=\int\eta(s_{\mathrm{dp}}\mid x)g(s_{\mathrm{dp}},\theta)\mathrm{d}s_{\mathrm{dp}}$ can be approximated with standard Monte Carlo integration techniques, as popular DP mechanisms such as Laplace and Gaussian mechanisms can be easily simulated. Using $M$ i.i.d. samples, the root-mean-squared-error (RMSE) to estimate $I(\theta,x)$ approximations are typically on the order of $\mathcal{O}(M^{-1/2})$ due to the central limit theorem.
In many applications, the DP query result $s_{\mathrm{dp}}$ serves as a private descriptive statistic of a dataset $x$. This statistic is commonly embedded in a low-dimensional space $\mathbb{S}$, where the dimension $r=\mathrm{dim}(\mathbb{S})$ is significantly less than the dimension of the data space $\mathrm{dim}(\mathbb{X}^{n})$. For the privacy mechanisms, we typically model the generation process as $\tau:(u,s(x))\mapsto s_{\mathrm{dp}}$ where $u\sim\mathcal{U}[0,1]^{r}$ is a uniform random variable from the $r$-dimensional hypercube. For example, the process $\tau(u,s(x))=s(x)-\frac{\Delta_{1}(s)}{\epsilon}\mathrm{sgn}(u-\frac{1}{2})\log\left[1-2|u-\frac{1}{2}|\right]$ achieves $\epsilon-$DP for 1-dimensional queries.
Randomized quasi-Monte Carlo (RQMC) methods, as detailed in (Owen, 1997a, ; Owen, 1997b, ), differ from traditional Monte Carlo (MC) methods in that it generate correlated, low-discrepancy sequences $\{v^{(1)},\cdots,v^{(M)}\}\subset[0,1]^{r}$. These sequences cover the parameter space more evenly than the pseudo-random sequences used by MC methods. This low-discrepancy feature helps to reduce the variance of the estimators, which is particularly useful in low-dimensional integration tasks (L’Ecuyer,, 2018). The estimator used in RQMC can be described as
@@
The RQMC estimator is unbiased and has a smaller RMSE compared to MC estimators. The $\star$-discrepancy of the point set $\{v^{(1:M)}\}$, denoted by $D^{\star}(v^{(1:M)})$ is of the order $\mathcal{O}(M^{-1}(\log M)^{r})=\mathcal{O}(M^{-1+\delta})$ for a positive constant $\delta$. If our neural approximation family $\tilde{g}_{\theta,x}(\cdot)$ has bounded Hardy-Krause variation $V_{\mathrm{HK}}[\tilde{g}_{\theta,x}(\cdot)]$ (Aistleitner et al.,, 2017), then, according to Basu and Owen, (2016), the mean squared error of the RQMC estimator in Equation (11) satisfies
$\mathbb{E}\left[\left(\hat{I}_{\theta,x}^{\mathrm{RQMC}}-I(\theta,x)\right)^{2}\right]\leq V^{2}_{\mathrm{HK}}(\tilde{g}_{\theta,x})(D^{\star})^{2}=\mathcal{O}(M^{-2+2\delta}).$ (12)
Thus, the RMSE of the RQMC estimator is of the order $\mathcal{O}(M^{-1+\delta})$, achieving faster convergence than an MC estimator with $\mathcal{O}(M^{-1/2})$. We verify this improvement in convergence from the RQMC estimator on the SIR model and linear regression example with the neural spline flow approximation family (Durkan et al.,, 2019), in Appendix E, Figure 7.
@@ 3.5 Sequential neural estimations
This section presents our sequential neural approximation methods on privacy-protected data. Our central goal is to approximate the private data posterior distribution $\pi(\theta\mid s_{\mathrm{dp}})$ given observed privatized data query $s_{\mathrm{dp}}$. We summarize the Sequential Private Posterior Estimation (SPPE) algorithm in Algo.1 and the Sequential Private Likelihood Estimation (SPLE) in Algo.2.
@@
     Update $\phi\leftarrow\arg\min_{\phi}\hat{\ell}_{\mathrm{PPE-A}}(\phi)$ using $\hat{I}^{\mathrm{RQMC}}$ Equation (11) on $\mathcal{D}_{r}$
@@
return $\hat{\pi}(\theta\mid s_{\mathrm{dp}}^{o})=q_{\phi}(\theta\mid s_{\mathrm{dp}}^{o})$
@@
     Update $\phi\leftarrow\mathop{\arg\min}_{\phi}\hat{\ell}_{\mathrm{PLE}}(\phi)$ using $\hat{I}^{\mathrm{RQMC}}$ Equation (11) on $\mathcal{D}_{r}$
@@
Both SPPE and SPLE use normalizing flows as the variational family to minimize some KL divergence, which takes the general form of Equation (9). We have designed their sequential training procedures to be sample efficient, in the sense that training data generated during previous rounds are kept and used in subsequent rounds.
In a sequential approximation procedure, we iteratively refine the neural approximations towards the target distribution. After the $r$-th training round, we incorporate the current neural density estimator $q_{\phi}$ into the proposal distribution of the next training round, using the automatic posterior transformation weights and loss functions described in Equations (7) and (8) respectively. Sequential training procedures can gradually move $q_{\phi}$ towards high-density regions of the private data posterior, and thus achieve good accuracy with fewer samples from the simulator, compared with non-sequential training procedures that use a fixed proposal distribution.
@@ 4.1 SIR model for disease spread
The SIR model is a time-series model that describes how an infectious disease spreads in a closed population. It is most often used as a deterministic ordinary differential equation (ODE), but can also be represented by a Markov jump process.
@@
Privatizing the infection curve.
Here, we propose a mechanism to privatize the infection trajectory $I(t)/K$, which is the proportion of infected individuals at each $t$.
Proposition 4 (DP infection trajectory).
@@
Consider a sequence of $L$ points $\{t_{1},\cdots,t_{L}\}$ in the time interval $[0,T]$, our privatized query can be $s_{\mathrm{dp}}=(s_{1},\cdots,s_{L})$ where each $s_{i}\sim\mathrm{Binomial}\left(n,\frac{I(t_{i})+m}{K+2m}\right)$ independently. The mechanism generating $s_{\mathrm{dp}}=(s_{1},\cdots,s_{L})$ satisfies $\epsilon$-DP, with $\epsilon=\frac{n}{m}L.$
This algorithm adds calibrated noise to the SIR process to produce $s_{\mathrm{dp}}$, a differential private time series. It can probably protect each individual’s infection status. We demonstrate that analysts can still make inferences about population parameters $(\beta,\gamma)$ by only knowing $s_{\mathrm{dp}}$, which retains information about the speed of disease spread.
Experiments on synthetic data.
@@
We illustrate the performance of SPPE and SPLE on synthetic privatized SIR model data. The data generating parameters are set to emulate a measles outbreak. We set $L=10$ time points with $n=1000$ and $m=1000$, achieving $\epsilon=10$ differential privacy. Lower privacy loss budgets ($\epsilon=1,0.1$) are explored in Appendix E.1.1 by adjusting $m$ accordingly. We also describe prior specifications and implementation details in the Appendix.
@@
Figure 1-A describes the convergence of the posterior approximations $\hat{\pi}(\theta\mid s_{\mathrm{dp}}^{o})$ towards the SMC-ABC (Beaumont et al.,, 2009) baseline, requiring up to $5\times 10^{5}$ simulations. We use SMC-ABC as the baseline because it does not resort to the variational approximations employed by SPPE and SPLE. Both SPPE and SPLE quickly adapt to meet the SMC-ABC results, with orders of magnitudes fewer simulations needed. After the first round of simulations, SPPE can identify the high probability region of $\pi(\theta\mid s_{\mathrm{dp}})$, while the SPLE posterior is still exploring the parameter space. See Appendix for the SPLE approximation after $r=1.$ After $r=5$ rounds, both methods can concentrate around the posterior mean and have captured the posterior correlation $\mathrm{cov}(\beta,\gamma\mid s_{\mathrm{dp}}).$ By inspecting marginal posterior histograms (Figure 6), we find that, in this example, the posterior approximated by SPPE is slightly more concentrated than those from SMC-ABC and SPLE.
@@
To quantitatively evaluate the performance of our methods, we use the following metrics: (1) MMD (Gretton et al.,, 2012): maximum mean discrepancy between the neural estimated posteriors and SMC-ABC posterior; (2) C2ST (Lopez-Paz and Oquab,, 2016): classifier two sample tests; (3) NLOG: negative log density at true parameters; and (4) LMD: log median distance from simulated to observed $s_{\mathrm{dp}}.$ Smaller values indicate better performance for all four metrics. In Figure 1-B, we compare the performance of these methods at various numbers of simulation samples/rounds. After Round 5, SPLE and SPPE have similar accuracy. We also compare their runtime in Table 1 of the Appendix. To achieve MMD lower than 0.1, SPPE is 6x faster and SPLE is 2x faster than SMC-ABC.
A
B
@@
Figure 1: Inference on SIR model. A. Convergence of sequential posterior estimations given DP-protected infection trajectory. Each round entails $N=1000$ simulations. B. Approximation accuracy by SPPE (orange) and SPLE (red) against the number of rounds, the error bars represent the mean with the upper and lower quartiles over 20 random trials.
@@
Figure 2: Inference on real infectious disease outbreaks. A. Visualization of the posterior distribution given private infection curve applied to flu, Ebola [in a) Guinea, b) Liberia, and c) Sierra Leone], and COVID-19 in Clark County, Nevada. All experiments use a privacy level of $\epsilon=10$. B. Mean and 95% credible intervals for $R_{0}=\beta/\gamma$ with different methods in each dataset. Grey: SMC-ABC; orange: SPPE; red: SPLE. A non-DP baseline (black stars) is also shown, which is obtained from the solution of the corresponding ordinary differential equations.
We apply our privacy mechanism and inference methods to several real infectious disease outbreaks: influenza, Ebola, and COVID-19. In Figure 2-A, we compare the posterior distributions $\pi(\beta,\gamma\mid s_{\mathrm{dp}})$ obtained from SMC-ABC, SPPE, and SPLE. The results are similar to synthetic data experiments: when SPPE and SPLE use the same computational resources, the SPPE posterior converges faster than SPLE. In Figure 2-B, we inspect the 95%-credible interval for the ratio $R_{0}=\beta/\gamma$, which is known as the basic reproduction number. Our $R_{0}$ estimates using privatized data are consistent with common estimates of $R_{0}$ for these diseases (Eisenberg,, 2020). Specifically, typical $R_{0}$ ranges are $(1.51,2.53)$ for Ebola, $(1.5,3.5)$ for COVID-19, and with the exception of the flu outbreak $(0.9,2.1)$, which should be modeled by the SI model instead of SIR.